OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)).

OpenStudy (zale101):

Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)). f(g(2)) f of g of 2 f(x^2-6x-7) but it's g of 2 not g of x so f((2)^2-(6(2)-7)=?

OpenStudy (zale101):

f((2)^2-(6(2)-7)=f(?)